# Rheological Analysis and Evaluation of Measurement Techniques for the Curing Polymethylmethacrylate Bone Cement in Vertebroplasty

### v1

@@ 2.3 Optical microscopy (tests Rh9 and Rh10)
Two additional tests were done on the rheometer to make bone cement samples for investigation with optical microscopy. In the first test (test Rh9), bone cement was prepared and simply left on the rheometer between the plates while maintaining 1.5 mm gap for 45 minutes without any action. In the second test (test Rh10), bone cement was prepared and subjected to 15 minutes of rotational 100 s shear rate, like in test Rh7b, and then 30 minutes of no action on the rheometer. The two cured bone cement discs formed as a result of these tests were investigated using an Axiotech microscope, which is a material microscope for non-transparent samples, equipped with a 2.5x objective. The images were captured using a digital camera Axiocam 105, and processed using software Axiovision 4.9.1. The microscope, camera, and processing software were provided by Carl Zeiss AG, Germany. Reflected light bright field technique was used for the illumination.
@@ 2.4 Tests on the custom-made injector (Tests Inj1 – Inj5)
The setup for the injection experiments is shown in Figure 0(a). The injector consisted of a carriage which was driven by a stepper motor using a ball screw with 5 mm feed per revolution. The stepper motor had a resolution of 1.8° corresponding to 200 steps per revolution, amounting to a calculated carriage resolution of 0.025 mm. The syringe was placed horizontally through a hole at the other end of the device as shown in the Figure 0(a) and held in place with a bolt. The cannula was then attached at the end of the syringe. The carriage pushed the plunger at the desired rate which could be programmed using a computer. At the point of pushing the plunger, a 200 N load cell was mounted on the carriage to measure the forces applied to the plunger during injection.
To prepare the bone cement, a 10 mL syringe was taken and sealed at its open end using a removable plug. The plunger of the syringe was removed and 3 mL of the monomer liquid was poured in it using a pipette. 7.8 grams of the PMMA powder were added directly into the syringe using a funnel and then mixed in the syringe using a PEEK stirring rod for 20 seconds (20 rotations). After mixing the components, the plunger was put back in the syringe and the plug at the other end was removed. The mixed bone cement was then first transferred to a 2 mL syringe until it was completely full, which was then used to fill up an 8-gauge cannula. The cannula had a side-opening and open front end, of which the side-opening was sealed using a strong adhesive tape so the bone cement could only come out of the front end when injected. The emptied volume of the 2 mL syringe was then again filled up using the remaining bone cement in the 10 mL syringe. At the beginning of the test, the carriage was set at the ‘unloaded’ position, which is far back to be able to mount the filled 2 mL syringe easily. Once the syringe and the cannula were mounted, the carriage was set to ‘loaded’ position, which was just behind the fully pulled back plunger. The ‘unloaded’ and ‘loaded’ positions were programmed in the device so it automatically arrived at them when commanded. At this point, the device was given the command to inject with a specified speed according to the required flow rate. The time from the start of mixing to the start of injection was measured using a stopwatch. The tests, named Inj1 to Inj5, were carried out with flow rates of 0.025, 0.05, 0.1, 0.2, and 0.4 mL/s respectively. During the injection, the load cell measured the force required for injection every 0.05 seconds. The time from mixing to start of injection in all tests was between 230 and 245 seconds.
@@ 2.5 Analytical calculations
Detailed derivations of these equations can be found in the works of 37, 38.
To apply these equations to our problem, the geometry of the apparatus was simplified as shown in Figure 0(b). The syringe and the cannula were assumed to be cylinders with the given dimensions. The coupling connecting the two parts had the same inner diameter as the cannula, hence its length was included in the cannula in the simplified geometry. The nozzle of the syringe was short in length and was therefore ignored. The change of cross-section from syringe to cannula would add to the pressure loss in reality, however this was ignored in our study. We then used the Equation 3 to obtain the pressure loss in our system. The pressure loss is then the sum of losses in the syringe and the cannula , given as
@@ 3.1 Oscillations at constant strain and frequency (Tests Rh1a and Rh1b)
At 0.2% strain amplitude and 23 °C, the complex modulus of the bone cement increased with time due to the ongoing polymerization. The polymerization occurred with a slow first phase, when the modulus increased gradually at a rate of about 29 kPa/min (Figure 1(a)). This is the working phase when the bone cement gradually hardens with time but maintains an injectable paste-like consistency. The slow first phase was followed by a rapid curing phase after about 20 minutes when the bone cement polymerized rapidly, and became solid in the next 5-10 minutes. The complex modulus after solidification was above 25 MPa. Interestingly, the loss factor always remained below 1, implying that the bone cement shows elastic prevalence, i.e. behaving more like an elastic solid than a liquid, even at the early stages of curing. The loss factor was between 0.5 and 0.7 until 25 minutes and rapidly decreased thereafter (Figure 1(a)).
@@
When the same test was carried out at higher 20% strain amplitude, the modulus was about two orders of magnitude lower (Figure 1(b)). Zooming in on the graph showed that initially, the modulus decreased with time. The decrease was an exponential decay at first and a continued decrease afterwards. The exponential decay is typical of viscoelastic stress relaxation behaviour, while the continued decrease afterwards could be probably because of fatigue resulting from repeated oscillations. After about 8 minutes, the modulus started to increase again, which was probably due to new polymer chains forming faster than the rate of breaking. The increase was, however, much slower compared to that observed in the test with 0.2% strain. Interestingly, the loss factor here was greater than 1, implying that the bone cement behaves more like a fluid when it is being significantly deformed. The loss factor did not go lower than 1 until after 25 minutes, when the rapid curing phase had already begun.
The properties of the bone cement continuously evolve because of the curing reaction. Hence, it is important to evaluate the so-called mutation time, which is defined as the time required for change in a given property by a factor equal to the Euler’s number . Taking the complex modulus as the property of interest, the bone cement has a mutation time of about 180 seconds and 700 seconds at 0.2% and 20% strain amplitude respectively. For reliable measurements, it is important that the measurement times are much less than mutation times to avoid the effect of the evolving properties in measurements. This is ensured to be the case for most of our measurements. As an exception, the measurement times are not negligible in the initial few points in the frequency sweep but they are still much smaller than mutation times.
@@
These results show that although the bone cement behaviour is predominantly solid when it is at rest or at very small deformations, while high deformations cause the bone cement to behave more like a fluid. This indicates that the bone cement is suitable for injections, despite this not being apparent from the test Rh1a. Furthermore, the curing process is inhibited by the deformations, as is evident from the much lower magnitude as well as a delayed and lower rate of increase in the complex modulus during the entire duration of the test.
@@ 3.2 Replication of injection (Tests Rh2a and Rh2b)
Figure 1(c) shows the results of the test Rh2a. During the initial stage of the test when the strain amplitude and frequency were low, the complex modulus increased gradually due to polymerization, slightly faster than at initial time of Rh1a (Figure 1(a)) test due to the lower strain used. In the next stage of the test, while it was subjected to higher strain and frequency, the modulus immediately dropped to 1% of the original magnitude and the loss factor jumped to a value higher than 1. This again showed the bone cement’s transition to more fluid-like behaviour as soon as it is deformed during polymerization. This was followed by an exponential decay of the modulus, as was also observed in the previous test Rh1b (Figure 1(b)). When the bone cement is brought to (almost) rest again in the final stage of the test (i.e. at about 12 minutes time), there is an immediate but small recovery of about 8% in the modulus. The polymerization is resumed and the rapid curing phase occurs at about similar time of 25 minutes, after which the bone cement solidifies. We note however that properties of bone cement modulated in this way (simulating more injection-like scenario), are different to those measured at low deformations in test Rh1a (Figure 1(a)). For example at the end of the test, the complex modulus is 2.5-fold lower and loss modulus is about 2-fold higher due to structural disturbances during curing. Carrying out the same test with the bone cement kept at 37 °C in the last stage of the test (Rh2b), there is hardly any recovery, however the rapid curing phase of the bone cement is triggered immediately. The bone cement is a fully cross-linked polymer solid at 22 minutes, as evident from the plateaued complex modulus and the loss factor value of 0.05.
The results show that the curing time is shortened when the bone cement is exposed to higher temperature. Once inside the vertebra of the patient, the bone cement starts curing rapidly. Vertebroplasty is often done in steps of smaller injections, in which case it is important for practitioners to be aware that the bone cement cures rapidly when exposed to the physiological body temperature. If the subsequent injections are not done in a timely manner, the viscosity of the cement inside the vertebra could get much higher than the one being injected. This viscosity difference could result in disadvantages like requiring higher injection force, unintuitive injection force development over the course of the injection, and unstable flow patterns inside the vertebra, as was highlighted in a previous work 19.
@@ 3.3 Step stress (Test Rh3) and step strain (Test Rh4)
Figure 2(a) shows the behaviour of the bone cement when subjected to step stress. The resultant strain was observed to be exponentially decaying, which is typical of viscoelastic behaviour, both when stress was applied and then removed. Upon removal of the stress, the strain did not converge back to zero, which indicated that there was permanent deformation in the sample. At very high stress, i.e. at 2000 Pa, the bone cement showed almost purely viscous behaviour as it deformed continuously, and there was no recovery after removal of the stress.
Figure 2(b) shows the stress response of the bone cement to step strain. Two tests were carried out with 1% and 100% step strain respectively. The resultant stress was maximum at the start of the applied step strain, and then decreased exponentially with time to a non-zero stress value, typical of viscoelastic stress relaxation behaviour. As the applied strain was removed, returning to zero strain state required stress in the opposite direction, which again showed similar relaxation behaviour. The decay curve could be fit to the relaxation equation
where is the time since the start of step strain, is the stress at time , is the stress at the start of relaxation, is the final stress the material relaxes to, and is the relaxation time. The relaxation times obtained from curve fitting are shown in Figure 2(b). Compared to the case of 1% strain, the relaxation times in the case of 100% step strain are lower, indicating that the bone cement relaxes faster owing to lesser resistance to deformation. This reinforces our previous observation that the bone cement has more fluid-like behaviour while it is being deformed.
@@ 3.4 Amplitude sweep (Test Rh5)
Figures 2(c) and 2(d) show the range and mean of the complex modulus, loss factor, and complex viscosity for three amplitude sweep tests on the bone cement. The modulus grew initially, likely because the strain was too small to inhibit the polymerization. At 0.02 - 0.04% strain, the modulus stopped growing and started lowering afterwards. The stress at which the complex modulus and the complex viscosity start to lower is the yield stress, which was observed to be around 70-110 Pa for this bone cement. This is the stress at which the bone cement starts to flow. Hence, this parameter is important to understand the injectability of the bone cement. The given yield stress range corresponds to injection force of about 0.7 N according to the equation obtained from Equation 2, which is relatively small and hence corroborates the use of bone cement in clinical injections. However, it is important to note that this value holds only for 295 20 seconds time since mixing, and would increase the more it is allowed to rest. In the 1-10% strain amplitude range, the loss factor went above 1. This is called the crossover point, or the flow point, and corresponded to about 500-1000 Pa stress ( 3-7 N of injection force). Finally, there was a sudden drop in the modulus near 200% strain or 2500-4000 Pa stress mark, which was likely because of insufficient friction between the rheometer plates and the bone cement, causing the plates to slip. The evidence for this phenomenon, also called wall slip, is provided in later sections.
@@ 3.5 Frequency sweep (Test Rh6)
Figures 2(e) and 2(f) show the range and mean of the complex modulus, loss factor, and the complex viscosity over three frequency sweep tests on the bone cement. We observed that the complex modulus kept increasing throughout the duration of the test. On the other hand, the complex viscosity initially increased slightly, then kept dropping as the angular frequency increased beyond 1.3 rad/s. This angular frequency can be considered the inception point of shear-thinning behaviour. The shear-thinning of the viscosity according to power law, i.e. Equation 1, started at about 2.6 rad/s. Fitting the curve to the power law equation yields the power law flow index . Value of between 0 and 1 indicates shear-thinning, which is expected for the PMMA bone cement. Measurements at higher angular frequency could not be executed due to the rheometer limitation of maximum frequency of 100 Hz.
@@ 3.6 Shear rate dependence (Tests Rh7a, Rh7b, and Rh8)
Figure 3(a) shows the results from the various steady shear rate sweep tests performed in rotational mode, also known as flow curves. The viscosity values were not affected by whether the test was done over a shorter shear rate ranges or all the way from 10 to 10 s. This implied that the duration of the test did not affect the results, at least within the first 10 minutes of mixing. The influence of curing was seen at very low shear rates in the order of 10 s, where the viscosity increased. The shear-thinning is steady until about 1 s. Upon curve-fitting, the flow index was obtained in the range of 0.20-0.30, and the consistency index in the range of 700-1100 Pa s.
@@
As the shear rate went above 1 s, the viscosity drop was noticeably steeper, until it started plateauing at shear rates greater than 20 s. The values of were less than zero in this range, implying that the shear stress reduced as the shear rate was increased, as seen in Figure 3(b). This could be a result of structural breakdown of the material or a measurement artefact due to the suspected wall slip effect. Performing the same tests using sandpaper attached to both the top and bottom plates (see test Rh7b in ‘Experimental’ section), the results showed a slight improvement in the maximum shear stress achieved (Figures 3(a) and 3(b)). This provided evidence that the drop in shear stress indeed occurred due to wall slip. This was confirmed from visual inspection through microscopy, as detailed in a later section.
@@
Figure 3(c) shows results of tests Rh8, i.e. time evolution at various constant shear rates. At the lowest shear rate 10 s, the bone cement showed a linear increase in viscosity with time of about 480 kPa s/min. The tests with higher shear rate showed lower magnitudes and slower increase with time. At 1 s shear rate, the viscosity kept decreasing for almost the entire duration of the test. At 100 s shear rate, the viscosity dropped almost entirely in the early stages (first 30 seconds) of the test . Viscosity values taken at various points in time for each of these tests were used to calculate the flow index with time. For this, the viscosity for shear rate 10 s was excluded since previous tests showed it was too low to induce shear-thinning. The results showed that the flow index was 0.33 in the beginning and 0.25 if the initial value from 100 s is taken into calculation. These values are in the same range as observed in tests Rh7a for shear rate less than 1 s. However, the value of kept dropping with time and became negative soon after the initial stages, as seen at 5 minutes mark in Figure 3(c). In fact, this was true even when the viscosity values from shear rates 1 and 100 s were excluded from the calculation for . This implied that the viscosity values only at the beginning of the test gave plausible shear-thinning characteristics. The same could be said for the consistency index, which when interpreted as the viscosity at unit shear rate, gave Pa s and kept decreasing with time.

### v2

@@ 2.3 Optical microscopy (tests Rh9 and Rh10)
Two additional tests were done on the rheometer to make bone cement samples for investigation with optical microscopy. In the first test (test Rh9), bone cement was prepared and simply left on the rheometer between the plates while maintaining a 1.5 mm gap for 45 minutes without any action. In the second test (test Rh10), bone cement was prepared and subjected to 15 minutes of rotational 100 s-1 shear rate, like in test Rh7b, and then 30 minutes of no action on the rheometer. The two cured bone cement discs formed as a result of these tests were investigated using an Axiotech microscope, which is a material microscope for non-transparent samples, equipped with a 2.5x objective. The images were captured using a digital camera Axiocam 105, and processed using software Axiovision 4.9.1. The microscope, camera, and processing software were provided by Carl Zeiss AG, Germany. Reflected light bright field technique was used for the illumination.
@@ 2.4 Tests on the custom-made injector (Tests Inj1 – Inj5)
The setup for the injection experiments is shown in Figure 1(a). The injector consisted of a carriage which was driven by a stepper motor using a ball screw with a 5 mm feed per revolution. The stepper motor had a resolution of 1.8° corresponding to 200 steps per revolution, amounting to a calculated carriage resolution of 0.025 mm. The syringe was placed horizontally through a hole at the other end of the device as shown in Figure 1(a) and held in place with a bolt. The cannula was then attached at the end of the syringe. The carriage pushed the plunger at the desired rate which could be programmed using a computer. At the point of pushing the plunger, a 200 N load cell was mounted on the carriage to measure the forces applied to the plunger during injection.
To prepare the bone cement, a 10 mL syringe was taken and sealed at its open end using a removable plug. The plunger of the syringe was removed and 3 mL of the monomer liquid was poured into it using a pipette. 7.8 grams of the PMMA powder were added directly into the syringe using a funnel and then mixed in the syringe using a PEEK stirring rod for 20 seconds (20 rotations). After mixing the components, the plunger was put back in the syringe and the plug at the other end was removed. The mixed bone cement was then first transferred to a 2 mL syringe until it was full, which was then used to fill up an 8-gauge cannula. The cannula had a side-opening and open front end, of which the side-opening was sealed using a strong adhesive tape so the bone cement could only come out of the front end when injected. The emptied volume of the 2 mL syringe was then again filled up using the remaining bone cement in the 10 mL syringe. At the beginning of the test, the carriage was set at the ‘unloaded’ position, which is far back to be able to mount the filled 2 mL syringe easily. Once the syringe and the cannula were mounted, the carriage was set to ‘loaded’ position, which was just behind the fully pulled back plunger. The ‘unloaded’ and ‘loaded’ positions were programmed in the device so it automatically arrived at them when commanded. At this point, the device was given the command to inject with a specified speed according to the required flow rate. The time from the start of mixing to the start of injection was measured using a stopwatch. The tests, named Inj1 to Inj5, were carried out with flow rates of 0.025, 0.05, 0.1, 0.2, and 0.4 mL/s respectively. During the injection, the load cell measured the force required for injection every 0.05 seconds. The time from mixing to the start of injection in all tests was between 230 and 245 seconds.
@@ 2.5 Analytical calculations
Detailed derivations of these equations can be found in the works of 43, 44.
To apply these equations to our problem, the geometry of the apparatus was simplified as shown in Figure 1(b). The syringe and the cannula were assumed to be cylinders with the given dimensions. The coupling connecting the two parts had the same inner diameter as the cannula, hence its length was included in the cannula in the simplified geometry. The nozzle of the syringe was short in length and was therefore ignored. The change of cross-section from syringe to cannula would add to the pressure loss in reality, however, this was ignored in our study. We then used the Equation 3 to obtain the pressure loss in our system. The pressure loss is then the sum of losses in the syringe and the cannula , given as
@@ 3.1 Oscillations at constant strain and frequency (Tests Rh1a and Rh1b)
At 0.2% strain amplitude and 23 °C, as shown in Figure 2(a), the complex modulus of the bone cement increased with time due to the ongoing polymerization. The polymerization occurred with a slow first phase, when the complex modulus increased gradually at a rate of about 29 kPa/min. This is the working phase when the bone cement gradually hardens with time but maintains an injectable paste-like consistency. The slow first phase was followed by a rapid curing phase after about 20 minutes when the bone cement polymerized rapidly, and became solid in the next 5-10 minutes. The complex modulus after solidification was above 25 MPa. The loss factor was about 0.5 - 0.7 until 25 minutes and rapidly decreased thereafter. Hence, the loss factor always remained below 1, implying that the bone cement was predominantly elastic, i.e. behaving more like an elastic solid than a liquid, even at the early stages of curing.
@@
When the same test was carried out at the higher 20% strain amplitude, the complex modulus was about two orders of magnitude lower, as shown in Figure 2(b). Zooming in on the graph showed that initially, the modulus decreased with time. The decrease was an exponential decay at first and a continued decrease afterwards. The exponential decay is typical of viscoelastic stress relaxation behavior, while the continued decrease afterwards could be probably because of fatigue resulting from repeated oscillations. After about 8 minutes, the modulus started to increase again, which was probably due to new polymer chains forming faster than the rate of breaking. The increase was, however, much slower compared to that observed in the test with 0.2% strain. Interestingly, the loss factor here was greater than 1, implying that the bone cement behaves predominantly like a fluid when it is being deformed, unlike when it is at rest. The loss factor did not go lower than 1 until after 25 minutes, which was deep in the rapid curing phase.
@@
These results show that the bone cement behavior is predominantly solid when it is at rest or with very small deformations, while high deformations cause the bone cement to behave more like a fluid. This indicates that the bone cement is suitable for injections, despite this not being apparent from the test Rh1a. Furthermore, the curing process is inhibited by the deformations, as is evident from the much lower magnitude as well as a delayed and lower rate of increase in the complex modulus during the entire duration of the test.
The properties of the bone cement continuously evolve because of the curing reaction. Hence, it is important to evaluate the so-called mutation time, which is defined as the time required for change in a given property by a factor equal to the Euler’s number 45. Taking the complex modulus as the property of interest, the bone cement was found to have a mutation time of about 180 seconds and 700 seconds at 0.2% and 20% strain amplitude respectively. For reliable measurements, the measurement times must be much less than mutation times to avoid the effect of the evolving properties in measurements. This is ensured to be the case for most of our measurements shown later in this work. As an exception, the measurement times are not negligible in the initial few points in the frequency sweep, but they are still much smaller than the mutation times.
@@ 3.2 Replication of injection (Tests Rh2a and Rh2b)
Figure 2(c) shows the results of the test Rh2a. During the initial stage of the test, i.e. when the strain amplitude and frequency were low, the complex modulus increased gradually due to polymerization. This rate was slightly faster than at the initial time of Rh1a (Figure 2(a)) test due to the lower strain used. In the next stage of the test, i.e. when it was subjected to higher strain and frequency, the modulus immediately dropped to 1% of the original magnitude and the loss factor jumped to a value higher than 1. This again showed the bone cement’s transition to more fluid-like behavior as soon as it is deformed during polymerization. This was followed by an exponential decay of the modulus, as was also observed in the previous test Rh1b (Figure 2(b)). When the bone cement is brought to (almost) rest again in the final stage of the test (i.e. at about 12 minutes), there is an immediate but small recovery of about 8% in the modulus. The polymerization is resumed and the rapid curing phase occurs at about a similar time of 25 minutes, after which the bone cement solidifies. We note, however, that the properties of bone cement modulated in this way (simulating a more injection-like scenario), are different to those measured at low deformations in test Rh1a (Figure 2(a)). For example at the end of the test, the complex modulus is 2.5-fold lower and the loss modulus is about 2-fold higher due to structural disturbances during curing. Carrying out the same test with the bone cement kept at 37 °C in the last stage of the test (Rh2b), there is hardly any recovery, however, the rapid curing phase of the bone cement is triggered immediately. The bone cement is a fully cross-linked polymer solid at 22 minutes, as evident from the plateaued complex modulus and the loss factor value of 0.05.
Figure 2(d) shows that the curing time is significantly shortened when the bone cement is exposed to higher temperature. Once inside the vertebra of the patient, the bone cement starts curing rapidly. A similar test was also done by Deusser et al. 29, in which they observed much slower curing, likely due to the torque-controlled nature of their measurement as opposed to our amplitude-controlled measurement. Nevertheless, they also observed a shift in the curing rate when the temperature was increased from 23°C to 37°C. Vertebroplasty is often done in steps of smaller injections, in which case practitioners need to be aware that the bone cement cures rapidly when exposed to the physiological body temperature. If the subsequent injections are not done in a timely manner, the viscosity of the cement inside the vertebra could get much higher than the one being injected. This viscosity difference could result in disadvantages like requiring higher injection force, unintuitive injection force development throughout the injection, and unstable flow patterns inside the vertebra, as was highlighted in a previous work 20.
@@ 3.3 Step stress (Test Rh3) and step strain (Test Rh4)
Figure 3(a) shows the behavior of the bone cement when subjected to step stress. The resultant strain was observed to be exponentially decaying, which is typical of viscoelastic behavior, both when stress was applied and then removed. Upon removal of the stress, the strain did not converge back to zero, which indicated that there was permanent deformation in the sample. At very high stress, i.e. at 2000 Pa, the bone cement showed almost purely viscous behavior as it deformed continuously, and there was no recovery after removal of the stress.
Figure 3(b) shows the stress response of the bone cement to step strain. Two tests were carried out with 1% and 100% step strain respectively. The resultant stress was maximum at the start of the applied step strain, and then decreased exponentially with time to a non-zero stress value, typical of viscoelastic stress relaxation behavior. As the applied strain was removed, returning to the zero strain state required stress in the opposite direction, which again showed similar relaxation behavior. The decay curve could be fit to the relaxation equation
where is the time since the start of step strain, is the stress at time , is the stress at the start of relaxation, is the final stress the material relaxes to, and is the relaxation time. The relaxation times obtained from curve fitting are shown in Figure 3(b). Compared to the case of 1% strain, the relaxation times in the case of 100% step strain are lower, indicating that the bone cement relaxes faster owing to lesser resistance to deformation. This reinforces our previous observation that the bone cement has a more fluid-like behavior while it is being deformed.
@@ 3.4 Amplitude sweep (Test Rh5)
Figures 3(c) and 3(d) show the range and mean of the complex modulus, loss factor, and complex viscosity for three amplitude sweep tests on the bone cement. The modulus grew initially, likely because the strain was too small to inhibit the polymerization. At 0.02 - 0.04% strain, the modulus stopped growing and started lowering afterwards. There are various definitions of the yield stress used in the literature. In this work, we call the stress at which the complex modulus and the complex viscosity start to lower as the yield stress, which was observed to be around 70-110 Pa for this bone cement. This is the point at which the ability of the bone cement to resist deformation starts to lower, and the bone cement begins to flow. Hence, this parameter is important to understand the injectability of the bone cement. The given yield stress range corresponds to injection force of about 0.83 N according to the equation obtained from Equation 2, which is relatively small and hence corroborates the use of bone cement in clinical injections. However, it is important to note that this value holds only for 295 20 seconds from the time of mixing, and would increase the more it is allowed to rest. Another interesting point is the crossover point, or the flow point, which is where the loss factor goes above 1, i.e. the bone cement transits to more fluid-like than solid-like behavior. This occurs in the 1-10% strain amplitude range or 500-1000 Pa stress ( 3-7 N of injection force). Finally, there was a sudden drop in the modulus near 200% strain or 2500-4000 Pa stress mark, which was likely because of insufficient friction between the rheometer plates and the bone cement, causing the plates to slip. The evidence for this phenomenon, also referred to as wall slip, is provided in later sections.
@@ 3.5 Frequency sweep (Test Rh6)
Figures 3(e) and 3(f) show the range and mean of the complex modulus, loss factor, and the complex viscosity over three frequency sweep tests done at 0.01% strain amplitude, i.e. within the linear viscoelastic limit as seen from the amplitude sweep test in Figure 3(c). We observed that the complex modulus kept increasing throughout the test. On the other hand, the complex viscosity initially increased slightly, then kept dropping as the angular frequency increased beyond 1.3 rad/s. This angular frequency can be considered the inception point of shear-thinning behavior. The shear-thinning of the viscosity according to power law, i.e. Equation 1, started at about 2.6 rad/s. Fitting the curve to the power law equation yields the power law flow index . A value of between 0 and 1 indicates shear-thinning, which is expected for the PMMA bone cement. Measurements at higher angular frequencies could not be executed due to the rheometer limitation of a maximum frequency of 100 Hz. The results of this test at other strain amplitudes are shown in upcoming sections for checking the validity of the Cox-Merz rule.
@@ 3.6 Shear rate dependence (Tests Rh7a, Rh7b, and Rh8)
Figure 4(a) shows the results from the various steady shear rate sweep tests performed in rotational mode, also known as flow curves. The measured viscosity was not affected by whether the test was done over smaller shear rate ranges or all the way from 10-3 to 103 s-1. This implied that the duration of the test did not affect the results, at least within the first 10 minutes of mixing. The influence of curing in the form of increasing viscosity was seen at very low shear rates in the order of 10-4 s-1. The shear-thinning is steady until about 1 s-1. Upon curve-fitting, the flow index was obtained in the range of 0.20-0.30, and the consistency index in the range of 700-1100 Pa sn.
@@
As the shear rate went above 1 s-1, the viscosity drop was noticeably steeper, until it started plateauing at shear rates greater than 20 s-1. The values of were less than zero in this range, implying that the shear stress reduced as the shear rate was increased, as seen in Figure 4(b). This could be a result of the structural breakdown of the material or a measurement artifact due to the suspected wall slip effect. Performing the same tests using sandpaper attached to both the top and bottom plates (see test Rh7b in ‘Experimental’ section), the results showed a slight improvement in the maximum shear stress achieved (Figures 4(a) and 4(b)). This provided first evidence that the drop in shear stress indeed occurred due to wall slip. This was confirmed from visual inspection through microscopy, as detailed in a later section.
@@
Figure 4(c) shows results of tests Rh8, i.e. time evolution at various constant shear rates. At the lowest shear rate 10-4 s-1, the bone cement showed a linear increase in viscosity with time of about 8 (kPa s)/s. At lower shear rates, i.e. at 10-3 s-1, 10-2 s-1, and 10-1 s-1, lower magnitudes and small dips in the viscosity can be seen, indicating retardation in the curing of the bone cement. This has also been observed for inorganic bone cements by Şahin and Kalyon 46. At 1 s-1 shear rate, the viscosity kept decreasing for almost the entire duration of the test. At 100 s-1 shear rate, the viscosity dropped almost entirely in the early stages (first 30 seconds) of the test. Viscosity values taken at specific points in time for each of these tests were used to calculate the flow index with time. For this, the viscosity for shear rate 10-4 s-1 was excluded since previous tests showed it was too low to induce shear-thinning. The results showed that the flow index was 0.33 in the beginning and 0.25 if the initial value from 100 s-1 is taken into calculation. These values are in the same range as observed in tests Rh7a for shear rate less than 1 s-1. However, the value of kept dropping with time and became negative soon after the initial stages, as seen at the 5-minute mark in Figure 4(c). In fact, this was true even when the viscosity values from shear rates 1 and 100 s-1 were excluded from the calculation for . This implied that the viscosity values only at the beginning of the test gave plausible shear-thinning characteristics. The same could be said for the consistency index, which when interpreted as the viscosity at unit shear rate, yields Pa sn in the beginning and then decreases with time. Note that the power law model gave similar fitting as the Herschel-Bulkley model, which is generally used for materials with yield stress. The comparison is provided in the supporting information in Figure S2 and Figure S3.